There are two MSP rings. Ring number one is conformed by NE_A, NE_B,NE_C and NE_D. Ring number two is conformed by NE_A,NE_F and NE_E.
The NE_A takes part of both rings. The MSP resources in NE_A are configured in the following protection groups:
ID # 2: 8-SLQ16-1 and 11-SLQ16-1
ID # 4: 8-SLQ16-3 and 11-SLQ16-3
The fiber resources of ring # 1 and ring # 2 are in the same cable. When they were broken, both rings switched normally. When the fiber was restored,
both rings entered to WTR status. After WTR time (10 min), the whole services of ring # 1 (protection group 2 in NE_A) became affected. The MSP protocol
in NE_A was in "Not started status".
Note: All ring members are OSN 3500
After fiber provider reported fiber restoration:
1. Confirm that there are no fiber problems: check optical power, fiber connection thorugh J0 byte and performance events.
2. Check the TU_AIS alarm to establish a common point in the services route.
3.Check the MS_APS_INDI_EX alarm parameters to determine the group ID that remains in switched state.
4. Verify the protocol status of the related rings using the protection subnet parameters.
5. If the protocol status is not started in any ring, stop and restart the protocol.
The MSP protection group 2 in was not restored in NE_A, meanwhile the NE_D protection was restored in time in ring # 1.
In accordance with the K events information, when fiber was fixed, NE_D sent 0x0120 , but the NE_A did not received the 0x0120 to switch idle page,
but the remote site (NE_D) switched to idle page normally, causing service affectation. The MSP protocol in NE_A for related protection group appeared
"Not started" for this reason.
(Please refer to the details in the attachement)
If there is a report that the fiber was retored, however the MS_APS_INDI_EX alarms is still in NM, check the alarms parameters to isolate the conflicting MSP gruop.
Be aware of the MSP handling process. Refer to the "Optix OSN Optical Transmission system Emergency Handling guide".